### v1

# Algebraic Connectivity Enhances Hyperedge Specificity in the Alzheimer’s Disease Continuum

Abstract
Functional MRI is a neuroimaging technique aiming at analyzing the functional activity of the brain by measuring blood-oxygen-level-dependent signals throughout the brain. The derived functional features can be used for investigating brain alterations in neurological and psychiatric disorders. In this work, we employed the hypergraph structure to model high-order functional relations across brain regions, introducing algebraic connectivity ($a(\mathcal{G})$) for estimating the hyperedge weights. The hypergraph structure was derived from healthy controls to build a common topological structure across individuals. The considered cohort included subjects covering the Alzheimer’s disease (AD) continuum, that is both mild cognitive impairment and AD patients. Statistical analysis was performed using the hyperedges’ weights as features to assess the differences across the three groups. Additionally, a mediation analysis was performed to evaluate the effectiveness and reliability of the $a(\mathcal{G})$ values, representing the functional information, as the mediator between tau-PET levels, a key biomarker of AD, and cognitive scores. The proposed approach outperformed state-of-the-art methods in identifying a larger number of hyperedges statistically different across groups. Among these, two hyperedges belonging to salience ventral attention and somatomotor networks showed a partial mediation effect between the tau biomarkers and cognitive decline. These results suggested that the $a(\mathcal{G})$ can be an effective approach for extracting the hyperedge weights, including important functional information that resides in the brain areas forming the hyperedges.
Keywords Alzheimer’s Disease  $\cdot$ Functional MRI  $\cdot$ Hypergraph  $\cdot$ Mediation Analysis  $\cdot$ Weights Computation
1 Introduction
Functional magnetic resonance imaging (fMRI) is a widely used neuroimaging technique for studying changes in blood-oxygenation-level-dependent (BOLD) signals in the brain. The BOLD signals reflect the activation or deactivation of the brain areas during time, thus allowing the study of synchronous and asynchronous regions’ temporal dynamics johnson2012brain . Functional connectivity (FC) measures can be derived starting from fMRI signals, typically computed as the Pearson correlation between the time series data of each pair of brain regions and represented as symmetric connectivity matrices. FC has been extensively employed in research for investigating brain functions and identifying functional alterations possibly associated with neurodegenerative and neuropsychiatric disorders. Usually, these studies employ FC matrices in conjunction with common statistical techniques or machine/deep learning approaches to extract meaningful information alorf2022multi ; geng2018multivariate ; plitt2015functional . From FC matrices, graph structures are extracted and used to investigate the network’s patterns. The graph is a data structure that, in the neuroscience field, can represent the brain, where each node of the graph represents a brain region, and the link between nodes is expressed as a measure of the association, typically the Pearson correlation gu2025fc ; cui2022braingb . Even if the graph-based structures and corresponding metrics and learning methods can better represent the brain functions and modulations, they still represent an approximation of a complex system due to the pairwise relationships expressed by those data structures.
This bottleneck can be naturally overcome by switching to hypergraphs. A hypergraph is a data structure consisting of hyperedges gathering multiple nodes. This allows capturing high-order relations across different brain regions, providing a more flexible representation with a higher number of degrees of freedom that is better suited for the problem at hand. By grouping different regions under the same hyperedge, complex relations across brain areas can be captured, relying on ad-hoc measures for hyperedge shaping. Some works have already applied functional hypergraphs in neuroscience studies, e.g., for disease classification and brain functional analysis teng2024constructing ; wang2024multiview ; liu2023deep ; gao2020hypergraph ; xiao2019multi .
However, many issues remain open, leaving room for further research. The two main challenges are i) building the hypergraph backbone structure, and ii) assigning the weights to the hyperedges. Regarding the first, different algorithms and techniques have been employed for detecting regions’ interplay and creating hyperedges composed of highly related regions starting from the fMRI time series. Previous works relied on the k-nearest neighbor algorithm to define the hyperedges and corresponding brain regions teng2024constructing ; zuo2021multimodal ; zhou2006learning . Other studies adopted clustering algorithms like k-means to learn the hypergraph structure ji2022fc ; zhou2006learning . The least absolute shrinkage and selection operator (LASSO) and Elastic Net with their related extensions are two other widely used methods that regress each time series with respect to all the others, hence extracting the functional relations across regions song2023brain ; yang2023constructing ; li2020hypernetwork ; xiao2019multi ; guo2018resting . The second challenge regards weight computation. The hyperedge weights are an essential part of the data structure since meaningful information can be extracted from those values, and subsequent analysis can be conducted starting from them. A common approach for assigning weights to the hyperedges is the Gaussian similarity kernel. A distance measure is computed between pairs of node features or time series, and then the sum or mean inside the hyperedge is calculated niu2023applications ; peng2023multiview ; gao2020hypergraph ; zhang2017joint . Another common approach to compute the hyperedge weights relies on the average or sum of the correlations across the regional time series of the same hyperedge bai2025hypergraph ; song2023brain ; shao2023analysis . Other approaches derive the weights using a learning procedure, such as optimization procedures or deep learning algorithms wang2024multiview ; ji2022fc ; cai2023discovering ; xiao2019multi .
In this work, we propose a novel method to compute hyperedge weights relying on the algebraic connectivity value computed over the brain regions data composing the hyperedges as the metric to define the hyperedge weights. The rationale for using the algebraic connectivity is due to the fact that it represents a measure of how well-connected the brain regions defining the hyperedges are. This metric allows us to compress in a unique value the information provided by the different brain areas, thus creating subgraph structures (hyperedges) which could describe differences in the functional activation across groups. To evaluate this method, we analyzed fMRI data from participants along the Alzheimer’s disease (AD) spectrum, including healthy controls (HC), individuals with mild cognitive impairment (MCI), and patients with AD. We performed a group-level statistical and mediation analysis on the hyperedge weights to assess differences in FC across groups and to assess the mediation effect of the functional data between tau-PET levels, a key biomarker of AD etiopathogenesis jack2016t , and cognitive decline.
In summary, the aims of this work are threefold: i) to derive hyperedge weights employing the algebraic connectivity measure in individuals across the Alzheimer’s disease continuum (HC, MCI, AD), based on a shared hypergraph structure; ii) to assess its effectiveness in differentiating diagnostic groups in relation to high-order FC patterns; and iii) to examine whether hyperedge weights derived from algebraic connectivity mediate the relationship between tau burden and cognitive decline.
2 Materials and Methods
2.1 Dataset
Neuroimaging rs-fMRI data used in the preparation of this article were obtained from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (https://adni.loni.usc.edu/). The ADNI was launched in 2003 as a public-private partnership, led by Principal Investigator Michael W. Weiner, MD. The primary goal of ADNI was to test whether serial MRI, PET, other biological markers, and clinical and neuropsychological assessment can be combined to measure the progression of MCI and early AD.
Table 1: Demographic information of the considered dataset.
Class Subjects Age Sex (M/F) ADAS-13 HC 310310310 73.1±8.273.1$\pm$ 8.273.1 ± 8.2 135/175 8.10±4.78.10$\pm$ 4.78.10 ± 4.7 MCI 199199199 73.8±11.073.8$\pm$ 11.073.8 ± 11.0 109/90 15.0±6.615.0$\pm$ 6.615.0 ± 6.6 AD 787878 76.7±8.176.7$\pm$ 8.176.7 ± 8.1 47/31 30.1±8.630.1$\pm$ 8.630.1 ± 8.6
A total of 587 subjects were selected and stratified along the AD continuum based on their clinical status, resulting in 310 HC, 199 MCI, and 78 AD. A subset of 217 HC subjects over the 310 HC was employed only for computing the hypergraph structure, named HC-hypergraph. In contrast, the remaining 93 HC (HC-analysis) individuals were used as a separate set only for the post-hoc analysis. Table 1 shows the demographic information of the dataset used in this work.
Rs-fMRI was acquired with the following acquisition protocol: TR/TE = 3000/30 ms, FA = 90°, FOV = 220 $\times$ 220 $\times$ 163 mm3, 3.4-mm isotropic voxel size. 200 rs-fMRI volumes were acquired in almost all subjects, with minimal variations in a small subset (e.g., 197 or 195 volumes). More details on data acquisition can be found in weiner2017alzheimer .
Standard preprocessing was applied to rs-fMRI data using FSL software (https://fsl.fmrib.ox.ac.uk/fsl/), followed by nuisance regression (6 motion parameters, white matter/CSF), pass-band filtering [0.01-0.08] Hz, scrubbing, and non-linear spatial normalization to the 2-mm MNI space DeBlasi2020 . Then, the mean regional time series of the Schaefer atlas yeo2011organization were extracted, hence obtaining 100 time series with 192 time points for each subject. Finally, each time series was z-scored, resulting in a zero mean and unitary standard deviation.
2.2 Hypergraph
The brain networks analysis can be performed through graph-based data structures, such as simple graphs and hypergraphs. A simple graph $\mathcal{G}$ is a data structure that relies on pairwise relations and can be represented as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{w})$, where $\mathcal{V}$ represents the set of vertices (nodes), $\mathcal{E}$ is the set of edges (connections between nodes), and $\mathbf{w}$ is the vector weights where each value is associated to each edge. Simple graphs have been widely used and investigated in brain functional analyses thanks to their ability in relating pairs of regions, observing synchronous and asynchronous activities and studying brain modulations and alterations in neurodegenerative and neuropsychiatric disorders stam2024hub ; stam2014modern . Despite their potential and widespread use, they have an important limitation when dealing with complex systems, like the human brain. In detail, the simple graphs can model only pairwise relations, thus limiting the understanding of the brain functionalities and behavior. To solve this limitation, the hypergraph, which is a complex data structure, can be suitable since each hyperedge that defines the hypergraph can connect more than two nodes. A hypergraph $\mathcal{HG}$ is defined as $\mathcal{HG}=(\mathcal{V},\mathcal{E},\mathbf{W})$, where the vertex set $\mathcal{V}=\{v_{1},v_{2},...,v_{N}\}$ is composed of N vertices, the hyperedge set =$\mathcal{E}=\{e_{1},e_{2},...,e_{M}\}$ is composed of M hyperedges, and $\mathbf{W}$ is the diagonal matrix of the hyperedge weights, where the $m$-th element in the diagonal corresponds to weight of the $m$-th hyperedge, $\mathbf{W}=diag(\{w(e_{1}),w(e_{2}),...w(e_{M})\}$. The hypergraph can be represented by an incidence matrix $\mathbf{H}\in\mathbb{R}^{N\times M}$, where
$\mathrm{H}_{n,m}=\left\{\begin{array}[]{l}1\quad\mathrm{if}\ v_{n}\in e_{m}\\ 0\quad\mathrm{otherwise}.\end{array}\right.$ (1)
Additionally, the incidence matrix $\mathbf{H}$ can also be non-binary, thus expressed as a weighted incidence matrix composed of real values (not only zeros and ones) in order to give a specific importance to each node in the different hyperedges. For each node, its degree represents the importance of the node in the $\mathcal{HG}$ across the hyperedges. Similarly, the hyperedge degree represents the number of nodes belonging to the hyperedge. This information in the hypergraph can be expressed in terms of diagonal matrices, the node and hyperedge degree diagonal matrices ($\mathbf{D}_{v}\in\mathbb{R}^{N\times N}$ and $\mathbf{D}_{e}\in\mathbb{R}^{M\times M}$, respectively) defined as follows:
[…]
where $d(v_{n})$ and $\delta(e_{m})$ represent the diagonal values for each node and hyperedge of $\mathbf{D}_{v}$ and $\mathbf{D}_{e}$, respectively. Finally, a hypergraph similarity matrix $\mathbf{S}\in\mathbb{R}^{N\times N}$ is defined as
$\mathbf{S}=\mathbf{HW}\mathbf{D}_{e}^{-1}\mathbf{H}^{T}$ (4)
where the elements of the matrix $\mathbf{S}$ represent high-order relations across the nodes of the hypergraph $\mathcal{HG}$ weighted by the hyperedge weights.
2.2.1 Hypergraph Structure Extraction
The $\mathcal{HG}$ structure in the form of the incidence matrix $\mathbf{H}$ was derived using LASSO regression, a well-established method for hypergraph structure computation jie2016hyper ; li2019multimodal ; xiao2019multi . LASSO regression was computed in order to find relations between the regional time series, using the $n$-th time series (region $n$) as the dependent variable and the other $N$-1 time series as the independent variables. For each time series $n$ in $1\leq n\leq N$, the optimization problem is defined as
$\underset{\boldsymbol{\alpha}_{n}}{\mathrm{min}}\frac{1}{2}||\mathbf{x}_{n}-\mathbf{X}_{n}\boldsymbol{\alpha}_{n}||_{2}^{2}\ +\ \lambda||\boldsymbol{\alpha}_{n}||_{1}$ (5)
$\mathrm{subject\ to}\quad\boldsymbol{\alpha}_{n}\succeq 0$ (6)
where $\mathbf{x}_{n}$ is the $n$-th time series used as the response variable, $\mathbf{X}_{n}\in\mathbb{R}^{P\times N-1}$ is the time series matrix of each subject, excluding the $n$-th time series (replaced by a zero vector), $\boldsymbol{\alpha}_{n}$ corresponds to the coefficients relating the response region with the other brain areas, and $\lambda$ is the sparsity hyperparameter used to extract a sparse representation of the $\boldsymbol{\alpha}_{n}$ vector. The $\lambda$ was optimized through 5-fold cross validation and set to 0.05.
The LASSO regression coefficients were extracted for all the centroid regions ($n-th$ region), resulting in a coefficient matrix $\mathbf{A}_{s}\in\mathbb{R}^{N\times N}$ for each subject $s$. Then, the distribution of the values in each $\mathbf{A}_{s}$ was computed and subsequently thresholded, considering the fifth percentile in order to remove the smallest and non-relevant coefficients. As further step, all $\mathbf{A}_{s}$ were binarized. After that, only the hyperedges with at least three brain regions, including the centroid region, were considered for the subsequent analysis, thus obtaining an incidence matrix $\mathbf{H}_{s}\in\mathbb{R}^{N\times M}$ for each subject.
Finally, to extract a common $\mathcal{HG}$ structure for all the individuals, the majority voting value of each cell of the incidence matrices $\mathbf{H}_{s}$ was computed, resulting in $\mathbf{H}$ that represented the $\mathcal{HG}$ structure in the form of incidence matrix, as described in Eq. (1). Figure 1 shows the pipeline used to extract the hypergraph structure.
Figure 1: Pipeline of this work. Initially, the mean time series are extracted from the fMRI volumes for each region of the Schaefer atlas. Then, the LASSO regression is computed between the intra-subject time series to extract the coefficients’ matrix, subsequently binarized in the incidence matrix $\mathbf{H}$. After that, the hyperedge weights (matrix $\mathbf{W}$) are computed. Finally, a post-hoc analysis is performed.
2.2.2 Hyperedge Modeling
Different methods have been proposed for computing the hyperedge weights ($\mathbf{W}$ matrix). The $\mathbf{W}$ matrix is a fundamental aspect when working with hypergraphs, since via this information different features can be extracted, such as the node degree matrix $\mathbf{D}_{v}$ and the hypergraph similarity matrix $\mathbf{S}$. To this end, in this work, we proposed a novel method of computing the weight of each hyperedge, relying on the algebraic connectivity ($a(\mathcal{G})$, Fiedler value) fiedler1973algebraic . The importance of $a(\mathcal{G})$ resides in the fact that it is a measure of how well connected a graph is de2007old . The $a(\mathcal{G})$ is computed on the subgraph composed of the regions belonging to a specific hyperedge. In detail, the absolute value of the FC matrix defined by the regional time series belonging to the hyperedge was computed, thus generating a subgraph of the brain regions of the hyperedge. The $a(\mathcal{G})$ is defined as the second-smallest eigenvalue extracted from the Laplacian matrix of the subgraph. Due to the limited number of regions involved in each hyperedge (at least three), LU Factorization was used to decompose the Laplacian matrix.
2.3 Post-hoc Analysis
2.3.1 Baseline Approaches
The $a(\mathcal{G})$ approach was compared via a post-hoc analysis with three common approaches to compute the hyperedge weights, and with the mean correlation value of the Schaefer functional networks (FNs), where each FN was considered as a hyperedge. In detail:
Gaussian similarity kernel
The Gaussian similarity kernel, defined as $K[i,j]=\exp{\frac{-||\mathbf{x}_{i}-\mathbf{x}_{j}||_{2}^{2}}{2\sigma^{2}}}$, was computed across the time series of the regions belonging to the hyperedge, where $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ are the $i$-th and $j$-th time series, respectively, and $\sigma$ was set using the heuristic median technique.
Mean Pearson correlation
The mean correlation was computed across all the pairwise correlations calculated between pairs of time series belonging to the hyperedge.
L2 norm of LASSO coefficients
The L2 norm of the LASSO coefficients in the hyperedge in the form of $\mathbf{w}_{m}=||\boldsymbol{\alpha}_{m}||_{2}^{2}$.
Mean FN correlation
The mean correlation values calculated from the seven FNs of the Schaefer atlas were employed as the hyperedge weights, where the hyperedges were expressed in terms of FNs regions.
2.3.2 Statistical Analysis
A statistical analysis was performed on the HC-analysis, MCI, and AD individuals, relying on the weights calculated for all the hyperedges. In detail, for each hyperedge, the non-parametric Kruskal-Wallis test was performed to identify the hyperedges that were statistically different across the three groups in terms of hyperedge weights. The non-parametric test was selected due to the non-normal distribution of the data assessed through the Shapiro-Wilk test.
Finally, a post-hoc analysis was performed to highlight the differences between the group pairs employing the Mann-Whitney test on the hyperedges that resulted in significant differences from the previous statistical test. The false discovery rate (FDR) strategy was used to correct the p-values for both analyses.
In order to check for any bias introduced using two different groups of HC for the hypergraph construction and for the post-hoc analysis, the same statistical analysis was also performed between the HC-hypergraph and HC-analysis individuals.
2.3.3 Mediation Analysis
Entorhinal cortex, a small region in medial temporal lobes, is among the earliest sites affected by AD pathology and is considered a key origin point for tau accumulation igarashi2023entorhinal . From this condition, pathological tau spreads to other brain regions via neural connections, contributing to disease progression and cognitive decline. The spread of tau pathology in AD is believed to follow a prion-like mechanism, advancing transneuronally across synaptically and functionally connected regions wang2024characterization . At the same time, tau deposition is known to disrupt FC, suggesting a reciprocal relationship between network integrity and tau progression roemer2025amyloid ; franzmeier2022tau . Motivated by this interplay, we conducted a mediation analysis to explore whether high-order functional features captured by hypergraph-derived hyperedge weights might mediate the link between entorhinal cortex tau burden, as measured by SUVR from tau-PET, and cognitive performance. The same analysis was performed on the hyperedge weights extracted from the Schaefer FNs. The analysis was conducted on a subset of the original cohort, including only participants with available entorhinal cortex tau SUVR values provided by ADNI and corresponding cognitive scores, and considering all diagnostic groups jointly. As a result, five MCI and three AD patients were excluded due to missing data. The following linear regressions modeled the relations between variables:
$\mathbf{y}$ $=\mathbf{i}_{1}+\mathbf{cz}+\mathbf{e}_{1}$ (7)
$\mathbf{w}_{m}$ $=\mathbf{i}_{2}+\mathbf{az}+\mathbf{e}_{2}$ (8)
$\mathbf{y}$ $=\mathbf{i}_{3}+\mathbf{c}^{\prime}\mathbf{z}+\mathbf{bw}_{m}+\mathbf{e}_{3}$ (9)
where $\mathbf{y}$ represents the cognitive composite scores (memory, executive, language, visual-spatial, and ADAS-13), $\mathbf{z}$ the tau SUVR in entorhinal cortex, $\mathbf{w}_{m}$ is the mediator variable, corresponding to the weight of the $m$-th hyperedge under analysis (and the FNs weights). $\mathbf{i}_{1,2,3}$, $\mathbf{e}_{1,2,3}$, and $\mathbf{c,a,b,c}^{\prime}$ are the intercepts, residual errors, and coefficients of the regression models, respectively. The $\mathbf{c}^{\prime}$ and $\mathbf{ab}$ define the direct and indirect effects of the mediation analysis, respectively, where the direct effect represents the effect of tau on the cognitive score without mediator (functional information), while the indirect one is the effect of tau SUVR on cognitive scores through the functional information (mediator). Figure 2 shows the graphical representation of the mediation analysis.
Figure 2: Schematic representation of the mediation analysis of this work.
Figure 3: Hyperedges found statistically relevant in the proposed analysis with their FDR-corrected p-values for the four different approaches.
Figure 4: Boxplots representing the distribution of the weights of the three groups for the statistically significant hyperedges and FNs.
All the hyperedges’ weights extracted using the $a(\mathcal{G})$ and the baseline approaches that resulted in statistically significant differences across the three groups were employed for the mediation analysis. Additionally, a bootstrap procedure with 1000 repetitions was employed to assess the significance of the indirect effect of the mediation analysis.
3 Results
3.1 Group Differences in Extracted Weights
A non-parametric statistical analysis was performed on the hyperedge weights extracted from all the approaches to assess differences across HC, MCI, and AD subjects. Figure 3 shows the brain regions forming hyperedges that exhibited statistically significant differences across groups. The corresponding FDR-corrected p-values are also reported. Table 2 summarizes the number of regions and their associated FNs for each significant hyperedge. Figure 4 highlights the distribution of these hyperedge weights, stratified by group and weighting method. Results showed that hyperedges 9, 48, 60, and 75, which include regions from the somatomotor, visual, and default mode FNs, exhibited reduced $a(\mathcal{G})$ in AD compared to both HC and MCI. In contrast, hyperedges 22 and 77, composed of regions within the dorsal and ventral attention FNs, showed increased $a(\mathcal{G})$ in AD patients. Similarly, the other approaches highlighted the same patterns with higher values in HC than MCI and AD for the hyperedges 9 and 75.
The weights computed with the $a(\mathcal{G})$ demonstrated greater discriminative power compared to the other three approaches, identifying six hyperedges with significant differences, while the mean Pearson correlation, the L2 norm of the LASSO coefficients, and the Gaussian similarity kernel identified only two significant hyperedges. All four methods highlighted differences across groups for hyperedge 9, which comprises regions within the somatomotor network. Hyperedge 75, composed of four regions from the salience ventral attention network and one region from the somatomotor network, was identified as significant by three of the methods, including the $a(\mathcal{G})$. When examining the FNs hyperedge weights, three out of seven FNs were found to be statistically different: the somatomotor, salience ventral attention, and default mode networks.
Table 2: Information of the statistically significant hyperedges.
Hyperedge Brain regions Functional networks involved 9 3 Somatomotor 22 3 Dorsal Attention (2), Ventral Attention (1) 48 5 Visual (2), Default (3) 59 5 Somatomotor (4), Ventral Attention (1) 60 7 Somatomotor (5), Ventral Attention (2) 75 5 Somatomotor (1), Ventral Attention (4) 77 3 Dorsal Attention (2), Ventral Attention (1)
To further assess pairwise group differences, post-hoc comparisons were performed using the Mann-Whitney U test. Table 3 summarizes the results of this analysis, highlighting the number of hyperedges that exhibited significant differences between each pair of groups (HC vs MCI, HC vs AD, and MCI vs AD).
Table 3: Post-hoc analysis results showing the number of different hyperedges between group pairs.
HC vs AD HC vs MCI MCI vs AD Algebraic connectivity 6/6 0/6 6/6 Gaussian similarity kernel 2/2 0/2 2/2 Mean Pearson correlation 2/2 0/2 2/2 L2 norm LASSO coeff. 2/2 1/2 2/2 Functional networks 3/3 1/3 2/3
The post-hoc analysis revealed a consistent ability across all considered approaches to discriminate HC from AD and MCI from AD. On the other hand, no hyperedge weights were able to differentiate the HC and MCI groups, except for the hyperedge 9 identified by the LASSO-based method. Similarly, the somatomotor network of the Schaefer atlas allowed differentiating HC and MCI groups.
Finally, the control analysis performed between the HC-hypergraph and HC-analysis found no differences between the two groups for all the methods, except for the Gaussian similarity kernel approach, which identified one hyperedge with an FDR-corrected $p<0.05$.
3.2 Mediation Analysis Results
Table 4: Statistically significant results of the mediation analysis with the bootstrap method (significance set at $p<0.05$) examining the relationship across tau entorhinal cortex SUVR, algebraic connectivity values, and five cognitive composite scores. Values represent the coefficients of the regression forming the total, direct, and indirect effects.
Hyperedge Algebraic connectivity - $a(\mathcal{G})$ Memory Executive Language Visual-Spatial ADAS-13 60 Total: -0.479 Direct: -0.466 Indirect: -0.014 Total: -0.214 Direct: -0.200 Indirect: -0.014 Total: 6.246 Direct: 6.076 Indirect: 0.170 75 Total: -0.479 Direct: -0.459 Indirect: -0.021 Total: -0.303 Direct: -0.283 Indirect: -0.019 Total: -0.214 Direct: -0.195 Indirect: -0.019 Total: -0.164 Direct: -0.149 Indirect: -0.015 Total: 6.246 Direct: 5.950 Indirect: 0.296
Table 4 summarizes the results of the mediation analysis performed using $a(\mathcal{G})$. Among the six hyperedges that showed significant group differences, only two revealed significant mediation effects. Specifically, hyperedge 60 showed a partial mediation effect between tau and the memory, language, and ADAS-13 composite scores. Hyperedge 75 revealed a broader significant partial mediation effect, linking the tau accumulation with all the cognitive scores. Across the analysis, the regression coefficients showed a similar pattern in the two hyperedges. In detail, a negative relation was found between tau burden and the five composite cognitive scores, where higher tau values were related to cognitive decline, e.g., lower values in the composite cognitive scores; while a positive effect was shown for the ADAS-13 score, where higher values were related to the disease.
The Gaussian similarity kernel and the mean Pearson correlation found significant mediation analysis for the weights of hyperedge 75 and all the cognitive scores, while the L2 norm of the LASSO coefficients approach found only significant relations in hyperedge 9 and for the memory and ADAS-13 scores. No significant mediation effects were found when using the Schaefer FNs-derived weights as mediator.
4 Discussion
The hypergraph is a data structure able to model high-order relations across brain regions, thus better approximating and representing a more realistic nature of the brain. The key concepts of the hypergraph are related to the hyperedges and their weights, since they represent groups of highly related brain regions and the importance of the different hyperedges in the hypergraph, respectively.
Selecting the most appropriate method for computing the hyperedge weight is essential, since it carries the functional information that resides in the brain areas of that subgraph. Assigning constant values or setting all weights to ones can be too simplistic yang2021graphlshc ; bai2021hypergraph and should be avoided since no meaningful features can be extracted. Similarly, assigning the hyperedge degree as the weight pisarchik2025hypergraph results in a partial loss of the regionals’ functional data, keeping only the information related to the number of regions forming the hyperedges.
In this study, taking advantage of the brain functional data, more complex approaches have been considered and developed for computing the hyperedges’ weights. Specifically, we explored approaches such as the Gaussian similarity kernel niu2023applications ; peng2023multiview ; gao2020hypergraph ; zhang2017joint , the L2 norm of LASSO coefficients balogh2022hypergraph ; gao2020hypergraph ; pan2021characterization , the mean or sum of pairwise correlations values between time series of the regions forming each hyperedge bai2025hypergraph ; song2023brain ; shao2023analysis , and the mean correlation values of the regions defined by the Schaefer functional atlas. With these approaches, we identified the hyperedges 9 (somatomotor areas) and 75 (somatomotor and ventral attention regions) and three FNs (somatomotor, salience, ventral attention, and default mode) that showed significant differences in AD versus MCI and AD versus HC, highlighting hipoconnectivity in the patients’ groups with respect to the HC one. These findings were consistent across the baseline approaches, including the outcomes of the Schaefer FNs, and with previous works where hipoconnectivity was observed in AD wiesman2021somatosensory ; li2012attention . On the other hand, an hyperconnectivity effect was featured in the hyperedge 59 (somatomotor and ventral attention areas) found by the L2 norm-based method, where higher connectivity was found in AD.
Even if these approaches obtained relevant results concordant with previous works, they relied on averaging or summing pair-wise metrics, thus not featuring the real nature of the hypergraph, except for the L2 norm applied to the LASSO coefficients that aggregated the regression coefficients. On these bases, the advantage of using the $a(\mathcal{G})$ derives from the fact that featuring the Laplacian matrix of the subgraph (hyperedge) and its decomposition to the second-smallest eigenvalue, we can extract a measure of how well the hyperedges are connected, and avoiding the use of mean of sum operations which brings approximation in the final results. The proposed method identified a larger number of hyperedges (9, 22, 48, 60, 75, and 77, where 9 and 75 were in common with the other approaches) statistically different across groups. These hyperedges included regions of the somatomotor, dorsal and ventral attentional, default mode, and visual networks. Interestingly, some hyperedges (e.g., 48, 60, and 75) grouped regions belonging to different FNs, such as somatomotor and attentional, and visual and default mode, which are usually considered separately in standard atlases. This suggests that our approach with hypergraphs could capture more insights and functional organization in AD that conventional methods overlook. Most of the significant hyperedges showed reduced $a(\mathcal{G})$ in AD patients, consistent with the idea of a disease-related loss of functional integration pini2021breakdown . In contrast, hyperedges involving dorsal and ventral attentional networks showed increased connectivity in the AD patients, reflecting possible compensatory effects. Further, the functional information that resided in regions belonging to the somatomotor and ventral attention FNs (hyperedges 60 and 75) showed a partial mediation between entorhinal tau load and cognitive performance, indicating that tau pathology may impair cognition by disrupting the coordination of large-scale networks in AD iordan2024salience ; sperling2019impact ; bejanin2017tau ; mielke2017association ; jack2016t . A similar pattern was previously observed employing the FC networks related to episodic memory and executive function as mediators between the tau levels and the memory and executive scores. They found that a significant mediation effect of FC was present between tau and memory function, but not in the executive ziontz2024behaviorally . Their results partially agree with our findings where for the hyperedges 60 and 75, significant mediation effects were present for memory score. On the contrary, only in the hyperedge 75 was a mediation effect between tau and executive function observed due to FC. Taken together, the obtained results highlighted the plausibility and potential of the proposed approach for deriving hyperedge weights that can capture meaningful functional information in the AD continuum and help understanding the neurophysiological mechanisms underlying Alzheimer’s Disease.
4.1 Main Limitations
One possible limitation of this work could be related to the LASSO method employed for computing the hypergraph structure. Even if this method has been widely used and the objective of this work was not to define new strategies for extracting the hypergraph structure, more performant approaches than LASSO could be considered for defining the hypergraph structure. Another possible limitation could be related to the $a(\mathcal{G})$ computation. This approach requires the graph to be connected, otherwise, it will result in a null score, thus limiting this approach only to connected graphs.
5 Conclusion
In this work, we proposed an approach for computing the hyperedge weights in the hypergraph structure relying on the algebraic connectivity $a(\mathcal{G})$, which is a measure of how well the brain regions composing the hyperedges are connected, even if they reside in different FNs. The results achieved with the $a(\mathcal{G})$ method showed a greater number of significantly different hyperedges than common approaches. The combination of decreased connectivity in sensorimotor, attentional, and default mode networks, together with increased connectivity in attention systems, underscores the heterogeneous and dynamic nature of the disease. Our use of $a(\mathcal{G})$ and hypergraph models provides a new way to capture these changes, highlighting both functional disconnection and potential adaptations. Furthermore, the mediation effects involving tau burden reinforce the idea of FC as the mechanistic link between molecular pathology and clinical symptoms, offering a pathway through which disease processes affect cognition.
Acknowledgments
Data collection and sharing for this project was funded by the Alzheimer’s Disease Neuroimaging Initiative (ADNI) (National Institutes of Health Grant U01 AG024904) and DOD ADNI (Department of Defense award number W81XWH-12-2-0012). ADNI is funded by the National Institute on Aging, the National Institute of Biomedical Imaging and Bioengineering, and through generous contributions from the following: AbbVie, Alzheimer’s Association; Alzheimer’s Drug Discovery Foundation; Araclon Biotech; BioClinica, Inc.; Biogen; Bristol-Myers Squibb Company; CereSpir, Inc.; Cogstate; Eisai Inc.; Elan Pharmaceuticals, Inc.; Eli Lilly and Company; EuroImmun; F. Hoffmann-La Roche Ltd and its affiliated company Genentech, Inc.; Fujirebio; GE Healthcare; IXICO Ltd.; Janssen Alzheimer Immunotherapy Research & Development, LLC.; Johnson & Johnson Pharmaceutical Research & Development LLC.; Lumosity; Lundbeck; Merck & Co., Inc.; Meso Scale Diagnostics, LLC.; NeuroRx Research; Neurotrack Technologies; Novartis Pharmaceuticals Corporation; Pfizer Inc.; Piramal Imaging; Servier; Takeda Pharmaceutical Company; and Transition Therapeutics. The Canadian Institutes of Health Research is providing funds to support ADNI clinical sites in Canada. Private sector contributions are facilitated by the Foundation for the National Institutes of Health (www.fnih.org). The grantee organization is the Northern California Institute for Research and Education, and the study is coordinated by the Alzheimer’s Therapeutic Research Institute at the University of Southern California. ADNI data are disseminated by the Laboratory for Neuro Imaging at the University of Southern California.
This study was funded by the Ministero dell’Università e della Ricerca (Bando PRIN 2022, “AI4BRAVE: AI for modeling of the Brain-Heart Axis in aging” project-reference code 202292PHR2), by the Ministero dell’Istruzione e del Merito (MIUR D.M. 737/2021, “AI4Health: empowering neurosciences with eXplainable AI methods” project), and by FSE project ANTIAGING: “Artificial intelligence for healthy aging” (project-reference code B31J23000910002) as well as NIH grant # R01AG073949 and # 1R01AG090597, and NSF grant # 2112455.
Additional information
The data used in this work were collected by ADNI (https://adni.loni.usc.edu/), and they are publicly available after requested on the ADNI website.